OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 16 when x = 12. Enter the constant of variation in lowest terms. y = x --------------------------------------------------------------------------------

OpenStudy (anonymous):

for anyone who needs this y varies directly with variable x y = kx, where k is any constant y=16; x=12 16 = 12k k=16/12 = 4/3 (in lowest terms)